# Moment-based parameter inference with error guarantees for stochastic reaction networks

### v1

@@ Appendix A Derivation of the Moment Equations
Assuming that $\pi$ is a stationary distribution, the equation
$\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{\pi}\coloneqq\sum_{\bm{x}^{\prime}\in\mathcal{X}}\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}^{\prime},\bm{x})\bm{x}^{\bm{\alpha}}\pi(\bm{x}^{\prime})=0$ (40)
@@
$=\sum_{r=1}^{R}k_{r}\sum_{|\bm{l}_{m}|\leq d}A^{d}_{\bm{\alpha}}(r,m)\bm{x}^{\bm{l}_{m}},$
where the elements in the coefficient matrix $A^{d}_{\bm{\alpha}}\in\mathbb{R}^{R\times N_{d}}$ are defined as
$A^{d}_{\bm{\alpha}}(r,m)\coloneqq\sum_{\begin{subarray}{c}\bm{\gamma}:\ \bm{l}_{m}-d_{b}\leq|\bm{\gamma}|\leq|\bm{\alpha}|-1\end{subarray}}\binom{\bm{\alpha}}{\bm{\gamma}}b_{r,\bm{l}_{m}-\bm{\gamma}}\bm{v}_{r}^{\bm{\alpha}-\bm{\gamma}},$ (42)
if $|\bm{l}_{m}|\leq|\bm{\alpha}|+d_{b}-1$ and zero otherwise. Dividing (41) by $s(\bm{x})$ and taking expectations it follows that $\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{\pi}=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\bm{y}_{:d}=0$, which are the moment equations (12) in the main text.
Appendix B Additional Moment Matrices Constraints
@@
$\zeta^{d}_{J}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{(N_{d}\times R)}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ M_{d}^{s}(\bm{z}_{\sigma(j)})\succeq 0,&s=0,\ldots,S,\\ &j=1,\ldots,J\end{array}}\right\},$ (44)
@@ Appendix C Observation Error Model
The idea of introducing models of the errors or noise associated with measured data in a study or experiment has been extensively discussed and implemented in the literature (Xu et al., 2020; Tang et al., 2020, 2023). This section briefly discusses how our approach can be applied to the case in which we consider an observation error model in the experiments. We call the moments of the measured numbers of molecules the measured moments ($\bm{y}^{\prime}_{:d}$) in contrast with the true moments ($\bm{y}_{:d}$). The main idea is to rewrite the sets in terms of the measured moments using a linear transformation between the true moments and the measured ones.
If the observation model leads to a linear relationship between the moments in the form
$\bm{y}^{\prime}_{:d}=B_{d}^{-1}\bm{y}_{:d}+\bm{c}_{d}$ (48)
where the invertible matrix $B_{d}\in\mathbb{R}^{{N_{d}}\times{N_{d}}}$ may correspond to a multiplicative error and a vector $\bm{c}_{d}\in\mathbb{R}^{N_{d}}$ may correspond to an additive error, then the moment equations in (12) satisfied by the true moments can be reformulated as equations satisfied by the measured moments,
$\bm{k}^{\top}A_{\bm{\alpha}}\bm{y}_{:d}=\bm{k}^{\top}A_{\bm{\alpha}}B_{d}(\bm{y}^{\prime}_{:d}-\bm{c}_{d})=0.$ (49)
Defining $\bm{z}^{\prime}_{j}=k_{j}\bm{y}^{\prime}_{:d}$, we can write
$\bm{z}_{j}=k_{j}\bm{y}_{:d}=k_{j}B_{d}(\bm{y}^{\prime}_{:d}-\bm{c}_{d})=B_{d}\bm{z}^{\prime}_{j}-k_{j}B_{d}\bm{c}_{d}.$ (50)
Assuming the matrix $B_{d}$ and the vector $\bm{c}_{d}$ are known given the error model, one forms sets as described in previous sections to constrain the parameters:
$\zeta^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}^{\prime}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z^{\prime}\in\mathbb{R}_{+}^{(N_{d}\times R)}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}^{\prime}_{l}\leq\bm{z}^{\prime}_{j}\leq k_{j}\hat{\bm{y}}^{\prime}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}B_{d}(Z^{\prime}-\bm{c}_{d}\bm{k}^{\top}))=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ B_{d}(\bm{z}^{\prime}_{j}-k_{j}\bm{c}_{d})^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(B_{d}(\bm{y}^{\prime}_{:d}-\bm{c}_{d}))\succeq 0,&s=0,\ldots,S\\ Z^{\prime}C^{\top}=\bm{y}^{\prime}\bm{c}^{\top}\\ \end{array}}\right\},$ (51)
where $\hat{\bm{y}}^{\prime}_{l},\hat{\bm{y}}^{\prime}_{u}$ are element-wise intervals of $\bm{y}^{\prime}$ and $Z^{\prime}\coloneqq\bm{y}^{\prime}\bm{k}^{\top}$, and we use $\bm{y}$ and $\bm{y}^{\prime}$ to denote $\bm{y}_{:d}$ and $\bm{y}^{\prime}_{:d}$, respectively, for notational simplicity.

### v2

@@ Appendix A Derivation of the Moment Equations
Assuming that $P_{t}$ satisfies the chemical master equation, the equation
$\frac{\mathrm{d}}{\mathrm{d}t}\langle x^{\bm{\alpha}}\rangle_{P_{t}}=\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{P_{t}}\coloneqq\sum_{\bm{x}^{\prime}\in\mathcal{X}}\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}^{\prime},\bm{x})\bm{x}^{\bm{\alpha}}P_{t}(\bm{x}^{\prime})$ (51)
@@
$=\sum_{r=1}^{R}k_{r}\sum_{|\bm{l}|\leq d}A^{d}_{\bm{\alpha}}(r,\bm{l})\bm{x}^{\bm{l}},$
where the elements in the coefficient matrix $A^{d}_{\bm{\alpha}}\in\mathbb{R}^{R\times N_{d}}$ are defined as in (16). Dividing (52) by $h(\bm{x})$ and taking expectations it follows that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu^{\bm{\alpha}}=\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{P_{t}}=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\bm{y}_{:d},$
which are the moment equations (15) in the main text.
Appendix B Additional Moment Matrices Constraints at Steady State
@@
$\zeta^{d}_{J}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{N_{d}\times R}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ M_{d}^{s}(\bm{z}_{\sigma(j)})\succeq 0,&s=0,\ldots,S,\\ &j=1,\ldots,J\end{array}}\right\},$ (54)
@@ Appendix C Observation Error Model
This section examines how our approach can be applied to the case of observation errors. We only describe this approach for the steady-state case, but the approach can be applied similarly to the transient case.
Let us denote the moments of the measured numbers of molecules the measured moments ($\bm{y}^{\prime}_{:d^{\prime}}$) in contrast with the true moments ($\bm{y}_{:d}$). We assume an affine relationship between the true and measured moments:
$\bm{y}^{\prime}_{:d^{\prime}}=B_{d^{\prime},d}\bm{y}_{:d}+\bm{c}_{d^{\prime}}$ (58)
where $B_{d^{\prime},d}\in\mathbb{R}^{{N_{d}^{\prime}}\times{N_{d}}}$ and $\bm{c}_{d^{\prime}}\in\mathbb{R}^{N_{d^{\prime}}}$.
For example, consider a one-dimensional SRN with an independent additive Gaussian error such that
$X^{\prime}=X+\epsilon,\ \epsilon\sim\text{N}(0,\sigma^{2}).$ (59)
Using the binomial expansion, one can then express each observed moment in terms of the true moments and the moments of the noise term. For instance, we have
$\begin{pmatrix}E(X^{\prime 0})\\ E(X^{\prime 1})\\ E(X^{\prime 2})\end{pmatrix}=\begin{pmatrix}1&0&0\\ E(\epsilon)&1&0\\ E(\epsilon^{2})&2E(\epsilon)&1\end{pmatrix}\begin{pmatrix}E(X^{0})\\ E(X^{1})\\ E(X^{2})\end{pmatrix},$ (60)
and thus, the error model satisfies the linear form (58).
Another common observation error model is the binomial model (Tang et al., 2020), where each molecule is detected with probability $p$:
$X^{\prime}|X\sim\text{Binomial}(X,p).$ (61)
This model also satisfies the linear error model (58), and the resulting matrix $B_{d^{\prime},d}$ depends explicitly on $p$.
Defining $\bm{z}^{\prime}_{j}=k_{j}\bm{y}^{\prime}_{:d^{\prime}}$ and defining the matrix form as $Z^{\prime}\in\mathbb{R}_{+}^{N_{d^{\prime}}\times R}$, we can write
$Z^{\prime}=B_{d^{\prime},d}Z+C_{d^{\prime}},$ (62)
where the $j$-th column in $C_{d^{\prime}}\in\mathbb{R}_{+}^{N_{d^{\prime}}\times R}$ is $k_{j}\bm{c}_{d^{\prime}}$.
Under the assumption that the matrix $B_{d^{\prime},d}$ and the vector $\bm{c}_{d^{\prime}}$ are known or can be estimated given the error model, one can form sets as described in the previous sections to constrain the parameters:
$\zeta^{d^{\prime},d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}_{:d}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\mathbb{R}_{+}^{N_{d}\times R}\\ \bm{y^{\prime}}_{:d^{\prime}}\in{\mathbb{R}_{+}^{N_{d^{\prime}}}}\\ Z^{\prime}\in\mathbb{R}_{+}^{N_{d^{\prime}}\times R}\end{array}}{\begin{array}[]{|ll}\bm{y}^{\prime}_{:d^{\prime}}=B_{d^{\prime},d}\bm{y}_{:d}+\bm{c}_{d^{\prime}}\\ Z^{\prime}=B_{d^{\prime},d}Z+C_{d^{\prime}}\\ k_{j}\hat{\bm{y}}^{\prime}_{:d^{\prime},l}\leq\bm{z}^{\prime}_{j}\leq k_{j}\hat{\bm{y}}^{\prime}_{:d^{\prime},u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y}_{:d})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}_{:d}\bm{c}^{\top}\\ \end{array}}\right\},$ (63)
where $\hat{\bm{y}}^{\prime}_{:d^{\prime},l},\hat{\bm{y}}^{\prime}_{:d^{\prime},u}$ are element-wise intervals on the observed moments.
Note that if the distribution of the error model is unknown, the relevant moments or parameters in $B_{d^{\prime},d}$ and $\bm{c}_{d^{\prime}}$ multiply $Z$ and thus lead to a non-linear, non-convex optimisation problem.